A certain group of test subjects had pulse rates with a mean of 72.5 beats per minute and a standard deviation of 11.1 beats per minute. Use the range rule of thumb to identify the limits separating values that are significantly low or significantly high. Is a pulse rate of 64.7 beats per minute significantly low or significantly​ high?
Answer #1

Solution :

Given that ,

mean = = 72.5

standard deviation = = 11.1

significantly low = - 2 = 72.5 - 2 * 11.1 = 50.3

significantly high = + 2 = 72.5 + 2* 11.1 = 94.7

64.7 beats per minute significantly low